# Bank–Laine functions with preassigned number of zeros

### v2

@@ 3.4. Proof for the case $n=1$
We first prove Theorem 1.4 for the case $n=1$. We shall consider the three cases (I), (II) and (III) together. Below we begin to estimate the numbers of poles and zeros of $F$, respectively. For two positive quantities $S_{1}(r)$ and $S_{2}(r)$, below we use the notation $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$.
We first estimate the number of poles of $F$. Let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$. Note that $\mathcal{N}_{k}\sim lk$ as $k\to\infty$ by (3.31). By the construction of $(m_{k})$ together with the definition of $\chi_{m,n}$ in (3.18) it follows that
@@
Moreover, we see from the choice of $k_{i}$ in (3.6) or (3.7) in the proof of Lemma 3.1 that $\sum_{i=1}^{k}m_{i}\sim(2\pi)^{-1}(\log(2\pi k_{i}))^{2}$ as $k\to\infty$ when $\lambda_{1}=0$ and that $\sum_{i=1}^{k}m_{i}\sim(2\pi)^{\lambda_{1}-1}k^{\lambda_{1}}$ as $k\to\infty$ when $0<\lambda_{1}\leq 1$. Together with these two relations, the two double inequalities in (3.41) imply that $n(r,\infty,G)\asymp(\log r)^{2}$ when $\lambda_{1}=0$ and that $n(r,\infty,G)\asymp r^{\lambda_{1}}$ when $0<\lambda_{1}\leq 1$. Then the relations in (3.40) imply that $n(r,\infty,F)\asymp(\log r)^{2}$ when $\lambda_{1}=0$ and that $n(r,\infty,F)\asymp r^{\lambda_{1}}$ when $0<\lambda_{1}\leq 1$. Thus $\lambda(f_{1})=\lambda_{1}$ in all three cases.
The estimate for the number of zeros of $F$ is similar. Let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$. By the construction of the sequence $(n_{k})$ together with the definition of $\chi_{m,n}$ in (3.18), we also have
@@
as $|z|\to\infty$ outside a set $\Omega_{4}$ of finite linear measure. Then the arguments for determining the order of $E$ and $A(z)$ are similar as that in the proof of Theorem 1.3. The lemma on the logarithmic derivative now implies that $E=F/F^{\prime}$ satisfies
@@
(3.51) $\begin{split}N(r,0,E)\leq c_{2}r^{\lambda_{2}}\end{split}$
for some positive constant $c_{2}$ and all large $r$. By the estimate in (3.49) together with the estimates in (3.50) or (3.51), the first main theorem of Nevanlinna theory [9, Chapter 1] implies that $T(r,E)=O(r)$. Thus $\rho(E)\leq 1$. On the other hand, by the definitions of $\tilde{g}_{m,n}$ in (3.18) and $\tau$ in (3.40) we easily see that $\log|E(z)|\geq c_{3}|z|$ for some positive constant $c_{3}$ and all $z$ on the negative real axis, implying that $\rho(E)\geq 1$. Hence $\rho(E)=1$.
Now, since $E$ is of finite order, it follows by the relation (1.8) together with the the lemma on the logarithmic derivative that
@@
Thus the first main theorem implies that $\rho(A)\leq 1$. On the other hand, by the relation in (1.8) and the definitions of $\tilde{g}_{m,n}$ in (3.18) and $\tau$ in (3.40), we easily see that $\log|A(z)|\geq c_{4}|z|$ for some positive constant $c_{4}$ and all $z$ on the positive real axis. Thus we also have $\rho(A)\geq 1$. Hence $\rho(A)=1$. This proves Theorem 1.4 for the case $n=1$.
@@ 3.5. Proof for the case $n\geq 2$
In this subsection, we show how to extend the results in the case $n=1$ to the case $n\geq 2$ for any positive integer $n$ with the same process as in [5, Section 5]. For two numbers $\tilde{\lambda}_{1},\tilde{\lambda}_{2}\in[0,n]$ such that $\tilde{\lambda}_{1}\leq\tilde{\lambda}_{2}$, we let $\lambda_{1}=\tilde{\lambda}_{1}/n$ and $\lambda_{2}=\tilde{\lambda}_{2}/n$. Then, we shall choose two sequences $(m_{k})$ and $(n_{k})$ associated with $\lambda_{1}$ and $\lambda_{2}$ in the same way as in Subsection 3.2. Note that $m_{k}=n_{k}=0$, $k=1,2,3$, for the two sequences $(m_{k})$ and $(n_{k})$ in any case.
Let $\chi:\mathbb{C}\to\mathbb{C}$ be the affine mapping such that $\chi(z)=x/l+iy$ where $l=1$ when $\lambda_{1}\leq\lambda_{2}<1$ and $l=3$ when $\lambda_{1}<\lambda_{2}=1$ and $l=4$ when $\lambda_{1}=\lambda_{2}=1$. Then we choose the numbers $\rho_{0}$ and $\sigma_{0}$ in [5, Section 5] to be $\rho_{0}=\sigma_{0}=1$ and the quasiconformal mapping $Q$ there to be $Q(z)=\chi(z)$.
We define a similar function $G_{0}(z)$ as in previous section. Denote by $J^{+}$ the preimage of the half-strip =$=\{z\in\mathbb{C}:\text{Re}\ z\geq 0,0\leq\text{Im}\ z\leq 2\pi\}$ under $z\to\chi(z)$ and $J^{-}=\{z\in\mathbb{C}:\text{Re}\ z\leq 0,0\leq\text{Im}\ z\leq 2\pi\}$. With $s_{0}=\log\log 2$, we then have
(3.52) $G_{0}(z)=\left\{\begin{array}[]{ll}\exp\exp(\chi(z)+s_{0}),&\text{if}\quad z\in J^{+},\\ \exp\exp(z+s_{0}),&\text{if}\quad z\in J^{-}.\end{array}\right.$
We note that $J:=J^{+}\cup J^{-}$ is bounded by the real axis and a line in the upper half-plane which, as the real axis, is mapped to $(1,\infty)$ by $G_{0}$.
In addition to $G_{0}$, we will also consider a modification $G_{1}$ of $G_{0}$ defined as follows. Let $U$ and $V$ be as in Subsection 3.2. For $\text{Re}\ z\geq 0$ we define
(3.53) $U_{1}(z)=\left\{\begin{array}[]{lll}U(z-\pi i),&\text{if}\quad\text{Im}\ z\geq\pi,\\ \exp(-\exp(z+s_{0}),&\text{if}\quad-\pi<\text{Im}\ z<\pi,\\ U(z+\pi i),&\text{if}\quad\text{Im}\ z\leq-\pi\end{array}\right.$
@@
(3.54) $V_{1}(z)=\left\{\begin{array}[]{lll}V(z-\pi i),&\text{if}\quad\text{Im}\ z\geq\pi,\\ \exp(-\exp(z+s_{0}),&\text{if}\quad-\pi<\text{Im}\ z<\pi,\\ V(z+\pi i),&\text{if}\quad\text{Im}\ z\leq-\pi.\end{array}\right.$
@@
(3.55) $G_{1}(z)=\left\{\begin{array}[]{ll}\exp(-\exp(\chi(z)+s_{0})),&\text{if}\quad z\in J^{+},\\ \exp(-\exp(z+s_{0})),&\text{if}\quad z\in J^{-}.\end{array}\right.$
@@
Then the quasiregular function $G$ for the case $n\geq 2$ is defined by slightly gluing the functions $G_{0}$ and $G_{1}$ in the same manner as in [5, Section 5]. For $j=1,\cdots,2n$ we put
(3.56) $\begin{split}\Sigma_{j}=\left\{z\in\mathbb{C}:(j-1)\frac{\pi}{n}<\arg z<j\frac{\pi}{n}\right\}.\end{split}$
Here we use the notation $n$ instead of the $N$ in [4, Section 5]). Let $H=\{z\in\mathbb{C}:\text{Im}\ z>0\}$ be the upper half-plane and $L=H/J$. Denote by $\bar{L}$ the reflection of $L$ on the real axis; that is, $\bar{L}=\{z\in\mathbb{C}:\bar{z}\in L\}$. Denote by $D_{j}$ the preimage of $L$ or $\bar{L}$ under the map $z\to z^{n}$ in $\Sigma_{j}$ for $j=1,\cdots,2n$. We will define the quasiregular map $G:\mathbb{C}\to\mathbb{C}$ which satisfies $G(z)=G_{0}(z^{n})$ for $z\in D_{j}$ if $12\leq j\leq 2N-$ 1 and $G(z)=G_{1}(z^{n})$ for $z\in D_{1}$ and $z\in D_{2n}$. In the remaining part of the plane we will define $G$ by a suitable interpolation which will require only (3.52) and (3.55). In order to do this, we will first define $G$ in a neighborhood of $\infty$. Thus, for a suitable $r_{0}>0$, we have to define $G$ in certain neighborhoods of the rays $\{z\in\mathbb{C}:\arg z=j\pi/n,|z|>r_{0}\}$. We shall not repeat the details.
Finally, by the same arguments as there we obtain a locally univalent meromorphic function $F$ such that $G=F\circ\tau$ for a quasiconformal mapping $\tau$ in the plane such that $\tau(z)\sim z$ as $z\to\infty$. Then we can estimate the numbers of poles and zeros of $G_{0}$ or $G_{1}$ in the sectors $D_{j}$ defined above in similar way as in the case $n=1$. Moreover, by the definitions of $G$ in the remaining parts $\mathbb{C}-\sum_{j=1}^{2n}D_{j}$, we may estimate the modulus of the resulting Bank–Laine $E$ and also the corresponding coefficient $A(z)$ of the second order different equation and prove that $\rho(E)=\rho(A)=n$ in similar way as in the case $n=1$. We shall not repeat the details here. This completes the proof of Theorem 1.4.
@@ 4.1. Preliminary lemmas
We need to prepare some preliminary lemmas which are slight modifications of those in [18, Section 2]. We shall first provide the corrected versions of two lemmas in [18, Lemma 2.1 and Lemma 2.2]. Since the construction of the Bank–Laines $E$ in the case $\lambda(E)=1$ has been included in Theorem 1.4, here we only consider the case where $\lambda(E)>1$.
@@
    $\begin{split}m_{k}=\delta\gamma(\log(x+2\pi))^{-2}(2\pi k)^{\delta\gamma-1}+O\left((\log(x+2\pi))^{-2}k^{\delta\gamma-2}\right)+O(1),\quad k\to\infty.\end{split}$
@@
For two numbers $\gamma\in(1,\infty)$ and $\delta\in[0,1]$, we choose $(m_{k})$ to be the sequence constructed in Lemma 4.1. Note that $(m_{k})$ can be modified to have the properties in Lemma 3.1 in the case $\delta=0$. Then we may follow the proof in [18, Lemma 2.1] to find a suitable sequence $(n_{k})$ such that the sequence $(N_{k})$ where $N_{k}=m_{k}+2n_{k}+1$ has the desired properties in Lemma 4.2 for the functions $h$ and $g$ there. Moreover, the assumptions that $m_{k}=0$ for $k=1,2,3$ in Lemma 4.1 and that $n_{k}=0$ for $k=1,2,3$ in Lemma 4.2 can be satisfied by slightly modifying the original proofs in [18].
In the original version of Lemma 4.2 (i.e.,[18, Lemma 2.1]), we have chosen the slopes of $h$ in the interval $[2\pi(k-1),2\pi k]$ to be $12n_{k}+$ 1. However, this then yields the function $g$ defined by $h(g(x))+\mathfrak{h}(g(x))=x^{\gamma}$ there deviates from the positive real axis too far away in the sense that $g(x)-x$ are unbounded as $x\to\infty$. For such a function $g$, the corresponding function $Q$ in the definition of $G$ in [18, Subsection 2.2] may not be quasiconformal in the plane. Now, by Lemma 4.2, we have the relations in (4.2) and (4.3) such that $g(x)-x$, as well as $g^{\prime}(x)-1$, is bounded as $x\to\infty$. With the corrected versions of Lemma 4.1 and Lemma 4.2, the proof for Theorem 1.2 is complete.
@@
(4.5) $\begin{split}\check{g}_{m,n}=g_{m,n}.\end{split}$
@@
for an affine mapping $\bar{\chi}:\mathbb{C}\to\mathbb{C}$ such that either $\bar{\chi}_{m,n}(z)=(z+1)/2$ or $\bar{\chi}_{m,n}(z)=z$. We choose the sequence $(m_{k})$ and $(n_{k})$ constructed in Lemma 4.1 and Lemma 4.2 respectively. The form of $\bar{\chi}$ defined in (4.6) will depend on the pair $(m_{k},2n_{k})$ of two nonnegative integers $m_{k}$ and $n_{k}$. The assumptions that $m_{k}=n_{k}=0$ for $k=1,2,3$ make it convenient to extend the construction of the entire function $A(z)$ in the case $\rho(A)=\rho\in(1/2,1)$ to the case $\rho(A)=\rho\in(n/2,n)$ for any positive integer $n$ as in [5, Section 5].
@@
$\begin{split}c=\frac{1}{M}\log\left[\frac{\binom{m_{1}+2n_{1}}{m_{1}}}{\binom{m_{2}+2n_{2}}{m_{2}}}\frac{M!}{N!}\right]-\log 2\quad\text{and}\quad\delta=\frac{1}{2}\min\{1,K\}.\end{split}$
@@
For each pair of $(m,2n)$ of two nonnegative integers $m$ and $n$, we shall work with the function $\bar{g}_{m,n}(z+\bar{s}_{m,n})$ instead of the function $\bar{g}_{m,n}(z)$ directly. Now, for the function $\bar{g}_{m,n}(z)$, we may use (4.15) or (4.16) and slightly modify the proof of [18, Lemma 2.4] to prove the following
@@
The estimate in (4.12) implies that $C\to 1$ as $N\to\infty$. Below we begin to analyse the asymptotic behaviors of $\bar{\phi}_{k}(z)$ defined in (4.7). For the choice of $(m_{k})$ and $(n_{k})$ in the case (I) we know that $\bar{\phi}_{1}=\phi_{1}$ is the identity in the plane and the asymptotic behaviors of $\bar{\phi}_{2}$ has been given in (4.8), so below we only consider the case $k\geq 3$. Then, by slightly modifying the proof of [18, Lemma 2.5], we have the following
@@
Below we shall always assume that the terms $O(1)$ appearing are only dependent on the constant $C$, not on other variables. By the estimates in (4.15) or (4.16), we may slightly modify the proof of [18, Lemma 2.6] to obtain the following

### v3

@@ 3.4. Proof for the case $n=1$
We first prove Theorem 1.4 for the case $n=1$. We shall consider the three cases (I), (II) and (III) together. We will use similar arguments as in the proof of Theorem 1.3 to estimate the numbers of poles and zeros of $F$ respectively and also to determine the orders of $E$ and $A$. Also, for two positive quantities $S_{1}(r)$ and $S_{2}(r)$, we use the notation $S_{1}(r)\asymp S_{2}(r)$, provided that there are two positive constants $c_{1}$ and $c_{2}$ such that $c_{2}S_{2}(r)\leq S_{1}(r)\leq c_{1}S_{2}(r)$ for all large $r$.
We first estimate the number of poles of $F$. Let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$. Note that $\mathcal{N}_{k}\sim lk$ as $k\to\infty$ by (3.31). By the construction of $(m_{k})$ together with the definition of $\chi_{m,n}$ in (3.18), we have
@@
By the choice of $k_{i}$ in (3.6) or (3.7) in the proof of Lemma 3.1, we easily see that $\sum_{i=1}^{k}m_{i}\sim(2\pi)^{-1}(\log(2\pi k))^{2}$ as $k\to\infty$ when $\lambda_{1}=0$ and that $\sum_{i=1}^{k}m_{i}\sim(2\pi)^{\lambda_{1}-1}k^{\lambda_{1}}$ as $k\to\infty$ when $0<\lambda_{1}\leq 1$. Together with these two relations, the two double inequalities in (3.41) imply that $n(r,\infty,G)\asymp(\log r)^{2}$ when $\lambda_{1}=0$ and that $n(r,\infty,G)\asymp r^{\lambda_{1}}$ when $0<\lambda_{1}\leq 1$. Then the relations in (3.40) imply that $n(r,\infty,F)\asymp(\log r)^{2}$ when $\lambda_{1}=0$ and that $n(r,\infty,F)\asymp r^{\lambda_{1}}$ when $0<\lambda_{1}\leq 1$. Thus $\lambda(f_{1})=\lambda_{1}$ in all three cases.
The estimate for the number of zeros of $F$ is similar. Again, let $r>0$ and choose $k\in\mathbb{N}$ such that $2\pi\mathcal{N}_{k-1}<r\leq 2\pi\mathcal{N}_{k}$. By the construction of the sequence $(n_{k})$ together with the definition of $\chi_{m,n}$ in (3.18), we have
@@
as $|z|\to\infty$ outside a set $\Omega_{4}$ of finite linear measure. The lemma on the logarithmic derivative now implies that $E=F/F^{\prime}$ satisfies
@@
(3.51) $\begin{split}c_{3}r^{\lambda_{2}}\leq N(r,0,E)\leq c_{2}r^{\lambda_{2}}\end{split}$
for two positive constants $c_{2},c_{3}$ and all large $r$. By the estimate in (3.49) together with the estimates in (3.50) or (3.51), the first main theorem implies that $T(r,E)=T(r,1/E)+O(1)=O(r)$. Thus $\lambda(E)\leq\rho(E)\leq 1$. In particular, if $\lambda_{2}=1$, then the estimate in (3.51) shows that $\lambda(E)=\rho(E)=1$. If $\lambda_{2}<1$, then the classical Hadamard factorization theorem together with the lemma on the logarithmic derivative easily yields that $T(r,E^{\prime}/E)=o(1)T(r,E)$ and $T(r,E^{\prime\prime}/E)=o(1)T(r,E)$ where $r\to\infty$. Then the relation in (1.8) implies that $T(r,A)=T(r,1/E)+o(1)T(r,E)=T(r,E)+o(1)T(r,E)$. Since an independent estimate below which only uses the fact that $E$ is of finite order yields $\rho(A)=1$, we also have $\rho(E)=1$.
Now, since $E$ is of finite order, it follows by the relation (1.8) together with the lemma on the logarithmic derivative that
@@
Thus the first main theorem implies that $\rho(A)\leq 1$. On the other hand, we note that $g_{1,1}$ is analytic in $\hat{\mathbf{\Pi}}_{1}^{+}$ as well as in $\check{\mathbf{\Pi}}_{-1}^{+}$. We see from the definition of $\tilde{g}_{m,n}$ in (3.18) that $\chi_{1,1}(z)=x/l+iy$ in the right half-plane and commutes with the complex conjugation. Together with the function $\tau$ in (3.40), we have $|(\chi_{1,1}\circ\tau^{-1})(z)|\leq 2|z|/l$ for $z$ such that $|z|$ is large. Denote $\tilde{\chi}_{1,1}=\chi_{1,1}+s_{1,1}$. Then by the same arguments as in the proof of Theorem 1.3 we may choose a curve in $(\tau\circ\tilde{\chi}_{1,1}^{-1})(\hat{\mathbf{\Pi}}_{1}^{+}\cup\check{\mathbf{\Pi}}_{-1}^{+})$, say $\mathcal{L}^{+}$, such that $(\tilde{\chi}_{1,1}\circ\tau^{-1})(\mathcal{L}^{+})$ is the positive real axis and also that $\tau\circ\tilde{\chi}_{1,1}^{-1}$ satisfies $|(\tau\circ\tilde{\chi}_{1,1}^{-1})^{\prime}(z)|\leq l|2z|$ for all $z\in(\tilde{\chi}_{1,1}\circ\tau^{-1})(\mathcal{L}^{+})$ such that $|z|$ is large. Then, by the relation in (1.8) as well as the definitions of $u_{m,n}$ in Subsection 3.2 and the relation $F=G\circ\tau^{-1}$, we easily see that $\log|A|\geq c_{4}|z|$ for some positive constant $c_{4}$ and all $z\in\mathcal{L}^{+}$ such that $|z|$ is large, implying that $\rho(A)\geq 1$. Hence $\rho(A)=1$. This proves Theorem 1.4 for the case $n=1$.
@@ 3.5. Proof for the case $n\geq 2$
In this subsection, we show how to extend the results in the case $n=1$ to the case $n\geq 2$ for any positive integer $n$ with the same process as in [5, Section 5]. For two numbers $\tilde{\lambda}_{1},\tilde{\lambda}_{2}\in[0,n]$ such that $\Lambda_{1}\leq\Lambda_{2}$, we let $\lambda_{1}=\Lambda_{1}/n$ and $\lambda_{2}=\Lambda_{2}/n$. Then, we choose two sequences $(m_{k})$ and $(n_{k})$ associated with $\lambda_{1}$ and $\lambda_{2}$ in the same way as in Subsection 3.2. Note that $m_{k}=n_{k}=0$, $k=1,2,3$, for the two sequences $(m_{k})$ and $(n_{k})$ in any case.
Let $\chi_{1,1}:\mathbb{C}\to\mathbb{C}$ be such that $\chi_{1,1}(z)=x/l+iy$ when $x\geq 0$ and $\chi_{1,1}(z)=x+iy$ when $x\leq 0$. Here $l=1$ when $\lambda_{1}\leq\lambda_{2}<1$ and $l=3$ when $\lambda_{1}<\lambda_{2}=1$ and $l=4$ when $\lambda_{1}=\lambda_{2}=1$. Then we choose the numbers $\rho_{0}$ and $\sigma_{0}$ in [5, Section 5] to be $\rho_{0}=\sigma_{0}=1$ and the quasiconformal mapping $Q$ there to be $Q(z)=\chi_{1,1}(z)$.
We define a similar function $G_{0}$ as in previous section. Denote by $J^{+}$ and $J^{-}$ the preimages of the half-strips $\{z\in\mathbb{C}:\text{Re}\ z\geq 0,0\leq\text{Im}\ z\leq 2\pi\}$ and $\{z\in\mathbb{C}:\text{Re}\ z\leq 0,0\leq\text{Im}\ z\leq 2\pi\}$ under $z\mapsto\chi_{1,1}(z)$ respectively. With $s_{0}=\log\log 2$, we then have
(3.52) $G_{0}(z)=\left\{\begin{array}[]{ll}\exp\exp(\chi_{1,1}(z)+s_{0}),&\text{if}\quad z\in J^{+},\\ \exp\exp(z+s_{0}),&\text{if}\quad z\in J^{-}.\end{array}\right.$
We note that $J:=J^{+}\cup J^{-}$ is bounded by the real axis and a line in the upper half-plane which, as the real axis, is mapped to $(1,\infty)$ by $G_{0}$. In addition to $G_{0}$, we will also consider a modification $G_{1}$ of $G_{0}$ defined as follows. Let $U$ and $V$ be as in Subsection 3.2. For $\text{Re}\ z\geq 0$ we define
$U_{1}(z)=\left\{\begin{array}[]{lll}U(z-\pi i),&\text{if}\quad\text{Im}\ z\geq\pi,\\ \exp(-\exp(z+s_{0}),&\text{if}\quad-\pi<\text{Im}\ z<\pi,\\ U(z+\pi i),&\text{if}\quad\text{Im}\ z\leq-\pi\end{array}\right.$
@@
$V_{1}(z)=\left\{\begin{array}[]{lll}V(z-\pi i),&\text{if}\quad\text{Im}\ z\geq\pi,\\ \exp(-\exp(z+s_{0}),&\text{if}\quad-\pi<\text{Im}\ z<\pi,\\ V(z+\pi i),&\text{if}\quad\text{Im}\ z\leq-\pi.\end{array}\right.$
@@
(3.53) $G_{1}(z)=\left\{\begin{array}[]{ll}\exp(-\exp(\chi_{1,1}(z)+s_{0})),&\text{if}\quad z\in J^{+},\\ \exp(-\exp(z+s_{0})),&\text{if}\quad z\in J^{-}.\end{array}\right.$
@@
For $j=1,\cdots,2n$ we put
$\begin{split}\Sigma_{j}=\left\{z\in\mathbb{C}:(j-1)\frac{\pi}{n}<\arg z<j\frac{\pi}{n}\right\}.\end{split}$
Here we have used the notation $n$ instead of the $N$ in [4, Section 5]). Let $H=\{z\in\mathbb{C}:\text{Im}\ z>0\}$ be the upper half-plane and $L=H/J$. Denote by $\bar{L}$ the reflection of $L$ on the real axis; that is, $\bar{L}=\{z\in\mathbb{C}:\bar{z}\in L\}$. Denote by $D_{j}$ the preimage of $L$ or $\bar{L}$ under the map $z\mapsto z^{n}$ in $\Sigma_{j}$ for $j=1,\cdots,2n$. We will define the quasiregular map $G:\mathbb{C}\to\mathbb{C}$ which satisfies $G(z)=G_{0}(z^{n})$ for $z\in D_{j}$ if $12\leq j\leq 2N-$ 1 and $G(z)=G_{1}(z^{n})$ for $z\in D_{1}$ and $z\in D_{2n}$. In the remaining part of the plane we will define $G$ by a suitable interpolation which will require only (3.52) and (3.53). In order to do this, we will first define $G$ in a neighborhood of $\infty$. Thus, for a suitable $r_{0}>0$, we have to define $G$ in certain neighborhoods of the rays $\{z\in\mathbb{C}:\arg z=j\pi/n,|z|>r_{0}\}$. If $j\not=1$ and $j\not=2n-1$, then $G$ is defined by the same expression in the domains $D_{j}$ and $D_{j+1}$ adjacent to the ray. (Here we have put $D_{2n+1}=D_{1}$. In similar expressions the index $j$ will also be taken modulo $2n$.) For $j=1$ and $j=2n-1$ the interpolation argument is more elaborate. Finally, the definition of $G$ can be extended to the remains.
By following similar process as in [4, Section 5]) we may finally obtain a locally univalent meromorphic function $F=G\circ\tau^{-1}$ for a quasiconformal mapping $\tau$ in the plane satisfying $\tau(z)\sim z$ as $z\to\infty$. We shall not repeat the details. Then $E=F/F^{\prime}$ is a Bank–Laine function and $E=f_{1}f_{2}$ for two normalized solutions of the second order linear differential equation (1.1) with an entire function $A$.
As in the case $n=1$, we may estimate the numbers of poles or zeros of $G_{0}(z^{n})$ in the sectors $D_{2},D_{3},\cdots,D_{2n-1}$ and that of $G_{1}(z^{n})$ in the sectors $D_{1},D_{2n}$, respectively. Since $G$ has no poles or zeros in the region $\mathbb{C}-\sum_{j=1}^{2n}D_{j}$, this finally yields that $n(r,\infty,F)\asymp(\log r)^{2}$ when $\Lambda_{1}=0$ and that $n(r,\infty,F)\asymp r^{\Lambda_{1}}$ when $0<\Lambda_{1}\leq n$ and also that $n(r,0,F)\asymp(\log r)^{2}$ when $\Lambda_{2}=0$ and that $n(r,0,F)\asymp r^{\Lambda_{2}}$ when $0<\Lambda_{2}\leq n$. Moreover, by the construction of $G$, we may estimate $\log|G|$ from above in a similar way as in the case $n=1$ to obtain that $\log|F|\leq\exp((1+o(1))|z|^{n}/l)$ as $z\to\infty$, which implies that $\rho(E)\leq n$ and also $\rho(A)\leq n$. On the other hand, by the definition of $G$ in the remaining parts $\mathbb{C}-\sum_{j=1}^{2n}D_{j}$, we may estimate $|A|$ from below along certain curve in similar way as in the case $n=1$. For example, letting $\eta:\mathbb{C}\to\mathbb{C}$ be the map such that $\eta(z)=z^{n}$, we may choose a curve in the closure of $(\tau\circ\eta^{-1}\circ\tilde{\chi}_{1,1}^{-1})(\eta^{-1}{(\Sigma_{2})}\cup\eta^{-1}(\Sigma_{3}))$, say $Y_{2}^{\prime}$, so that $(\tilde{\chi}_{1,1}\circ\eta\circ\tau^{-1})(Y_{2}^{\prime})$ is the positive real axis. For a point $z$ in the real axis, we may use the Cauchy formula as in the proof of Theorem 1.3 to show that $|(\tau\circ\eta^{-1}\circ\tilde{\chi}_{1,1}^{-1})^{\prime}(z)|\leq l|2z|^{1/n}$ when $|z|$ is large, which then yields that $\rho(A)\geq n$. Finally we conclude that $\rho(E)=\rho(A)=n$. We shall not repeat the details here. This completes the proof of Theorem 1.4.
@@ 4.1. Preliminary lemmas
We need to prepare some preliminary lemmas which are slight modifications of those in [19, Section 2]. We shall first provide the corrected versions of two lemmas in [19, Lemma 2.1 and Lemma 2.2]. Since the construction of the Bank–Laines $E$ in the case $\lambda(E)=1$ has been included in Theorem 1.4, here we only consider the case where $\lambda(E)>1$.
@@
    $\begin{split}m_{k}=\delta\gamma(\log(k+2\pi))^{-2}(2\pi k)^{\delta\gamma-1}+O\left((\log(k+2\pi))^{-2}k^{\delta\gamma-2}\right)+O(1),\quad k\to\infty.\end{split}$
@@
For two numbers $\gamma\in(1,\infty)$ and $\delta\in[0,1]$, we choose $(m_{k})$ to be the sequence constructed in Lemma 4.1. Note that $(m_{k})$ can be modified to have the properties in Lemma 3.1 in the case $\delta=0$. Then we may follow the proof in [19, Lemma 2.1] to find a suitable sequence $(n_{k})$ such that the sequence $(N_{k})$ where $N_{k}=m_{k}+2n_{k}+1$ has the desired properties in Lemma 4.2 for the functions $h$ and $g$ there. Moreover, the assumptions that $m_{k}=0$ for $k=1,2,3$ in Lemma 4.1 and that $n_{k}=0$ for $k=1,2,3$ in Lemma 4.2 can be satisfied by slightly modifying the original proofs in [19].
In the original version of Lemma 4.2 (i.e.,[19, Lemma 2.1]), we have chosen the slopes of $h$ in the interval $[2\pi(k-1),2\pi k]$ to be $12n_{k}+$ 1. However, this then yields the function $g$ defined by $h(g(x))+\mathfrak{h}(g(x))=x^{\gamma}$ there deviates from the positive real axis too far away in the sense that $g(x)-x$ are unbounded as $x\to\infty$. For such a function $g$, the corresponding function $Q$ in the definition of $G$ in [19, Subsection 2.2] may not be quasiconformal in the plane. Now, by Lemma 4.2, we have the relations in (4.2) and (4.3) such that $g(x)-x$, as well as $g^{\prime}(x)-1$, is bounded as $x\to\infty$. With the corrected versions of Lemma 4.1 and Lemma 4.2, the proof for Theorem 1.2 is complete.
@@
(4.5) $\begin{split}\check{g}_{m,n}(z)=g_{m,n}(z).\end{split}$
@@
for an affine mapping $\bar{\chi}:\mathbb{C}\to\mathbb{C}$ such that either $\bar{\chi}_{m,n}(z)=(z+1)/2$ or $\bar{\chi}_{m,n}(z)=z$. We choose the sequence $(m_{k})$ and $(n_{k})$ constructed in Lemma 4.1 and Lemma 4.2 respectively. The form of $\bar{\chi}$ defined in (4.6) will depend on the pair $(m_{k},2n_{k})$ of two nonnegative integers $m_{k}$ and $n_{k}$. The assumptions that $m_{k}=n_{k}=0$ for $k=1,2,3$ make it convenient to extend the construction of the entire function $A$ in the case $\rho(A)=\rho\in(1/2,1)$ to the case $\rho(A)=\rho\in(n/2,n)$ for any positive integer $n$ as in [5, Section 5].
@@
$\begin{split}c=\frac{1}{N}\log\left[\frac{\binom{m_{1}+2n_{1}}{m_{1}}}{\binom{m_{2}+2n_{2}}{m_{2}}}\frac{N!}{M!}\right]-\log 2\quad\text{and}\quad\delta=\frac{1}{2}\min\{1,K\}.\end{split}$
@@
For each pair of $(m,2n)$ of two nonnegative integers $m$ and $n$, we shall work with the function $\bar{g}_{m,n}(z+\bar{s}_{m,n})$ instead of the function $\bar{g}_{m,n}(z)$ directly. Now, for the function $\bar{g}_{m,n}(z)$, we may use (4.15) or (4.16) and slightly modify the proof of [19, Lemma 2.4] to prove the following
@@
The estimate in (4.12) implies that $C\to 1$ as $N\to\infty$. Below we begin to analyse the asymptotic behaviors of $\bar{\phi}_{k}(z)$ defined in (4.7). For the choice of $(m_{k})$ and $(n_{k})$ in the case (I) we know that $\bar{\phi}_{1}=\phi_{1}$ is the identity in the plane and the asymptotic behaviors of $\bar{\phi}_{2}$ has been given in (4.8), so below we only consider the case $k\geq 3$. Then, by slightly modifying the proof of [19, Lemma 2.5], we have the following
@@
Below we shall always assume that the terms $O(1)$ appearing are only dependent on the constant $C$, not on other variables. By the estimates in (4.15) or (4.16), we may slightly modify the proof of [19, Lemma 2.6] to obtain the following